# New Permutation Decomposition Techniques For Efficient Homomorphic Permutation

### v5

@@ 6.1. Homomorphic Matrix Transposition
Table 3 provides HMT benchmarks under dimension $d=128$ and decomposition depths $l\leq\log{d}-1$ (requiring multiplication depth $l+1$). In this setup, the leftmost permutation matrix in the decomposition chain of $U^{t}$ has approximately $d/2^{l}$ non-zero diagonals, which is automatically implemented with the BSGS algorithm.
Compared to the original undecomposed version (Jiang et al., 2018), the ideal decomposition yields significant complexity reductions, even when the decomposition depth is shallow. This aligns with the reduction in rotations, whose number changes from $d_{1}+\frac{2d}{d_{1}}$ before decomposition to $2\ell+d_{1}+\frac{2d}{2^{\ell}d_{1}}$ after decomposition. Notably, with $d_{1}=16$, our algorithm demonstrates a speed increase up to 50.2% and a rotation key reduction up to 30.0%.
@@ 6.2. Homomorphic Matrix Multiplication
Table 4 compares our enhanced HMM in Section 4.2 to the original scheme (Rizomiliotis and Triakosia, 2022). Unlike our variable $d^{2}d^{\prime}$ matrix encoding space, the original $d^{3}$-encoding allows limited choices of dimensions. Thus we enable comparisons across a broader range of dimensions by using the conjugate-invariant ring based CKKS (Kim and Song, 2019), which shifts the ciphertext’s encoding capacity to $N$ real numbers (HE parameters remain unchanged).
In the table, parameter $m=N/d^{2}d^{\prime}$ denotes the maximum number of matrices that can be computed in parallel. For the original encoding setting $d^{\prime}=d$, our scheme achieves up to a 70% speedup and a 42% reduction in rotation keys compared to the original algorithm. This is attributed to the permutation decomposition and $1$-padding idea, which allow the complexity to grow logarithmically with $d$ and remain independent of multiplication depth overhead. Test results under different $d^{\prime}$ demonstrate the enhancement of our more flexible encoding space utilization method. We can trade encoding space for higher parallelity and lower amortized time per matrix multiplication.
6.3. Application to Oblivious Neural Network Inference
Oblivious inference refers to the task of outsourcing an ML model without revealing its parameters or inputs, wherein the model owner is not required to run the model locally (Jiang et al., 2018; Chen et al., 2019; Rizomiliotis et al., 2022). Jiang et al. (Jiang et al., 2018) proposed an efficient HE-based framework, E2DM, to realize this objective, predicting the MNIST dataset (Deng, 2012) over an encrypted neural network. The computational complexity is dominated by their proposed HMM, which, as discussed above, maintains the best-known amortized complexity. We adopt their service paradigm but replace the HMM module in their framework with our version. Our objective is to demonstrate the practicality of our HMM by achieving lower latency when redundant encoding capacity is available, particularly under small-batch inference scenarios.
Table 5 describes the employed neural network. We call it E2DM-lite, as it is derived by halving the number of convolutional channels and neurons in E2DM, but achieves an accuracy of 97.7% (only 0.4% lower than the original version). We present benchmarks of E2DM-lite in Table 6. When implemented with the original HMM by Jiang et al., E2DM-lite requires an inference batch size of $128$ for optimal performance. Any smaller batch yields no significant time reductions. In contrast, our HMM takes advantage of the encoding space left by small batches, achieving latency reductions of up to 7.9x while maintaining an amortized runtime that is only slightly higher. We provide implementation details in Appendix E.
Table 6 also presents benchmarks of other state-of-art HE-based oblivious neural network inference frameworks on the MNIST dataset. Note that although they use similar underlying HE schemes (Cheon et al., 2017, 2019) and network topologies, we cannot force them to work under identical parameters of HE and neural network, which have been highly customized for different computation strategies. Therefore, the benchmarks are obtained under the original parameters of these frameworks, except for E2DM, which only ensures 80-bit security. We have it implemented under the same HE setting as our scheme.
Jiang et al.’s HMM implementation in E2DM and E2DM-lite exhibits good amortized performance under high batch sizes, but introduces relatively high depth, making it more suitable for shallow, wide networks. The strategy designed by Chen et al.(Chen et al., 2019) (CDKS) employs their proposed multi-key homomorphic encryption to avoid introducing a third-party authority. It has a good latency, but the batch size is relatively small. POSEIDON(Sav et al., 2021) is a federated neural network learning framework based on threshold HE. It provides fast oblivious inference after the network is homomorphically trained among parties. However, clients outside the training system need to encrypt their inquiries with the joint public key of the training parties.
@@ 6.3. Application to Oblivious Neural Network Inference
Convolution Input: $28\times 28$, kernel:7×$\times 7$ 7, stride:$3$, channels:$2$
@@
    Original implementation under Jiang et al.’s HMM.
@@
$\log{n}$ Time (ms) Depth Time (ms) Depth Speed-up
$8$ 1612.21 4 950.74 3 ×$\times 1.69$
$9$ 1816.51 3 1183.75 3 ×$\times 1.53$
$10$ 1687.24 5 1239.73 4 ×$\times 1.36$
$11$ 1909.86 5 1415.83 5 ×$\times 1.35$
$12$ 2198.14 6 1545.78 6 ×$\times 1.42$
$13$ 1887.42 8 1632.65 7 ×$\times 1.15$
$14$ 2278.72 8 1797.25 8 ×$\times 1.27$
@@ 6.4. Arbitrary Homomorphic Permutation
Table 7 compares our proposed homomorphic permutation algorithm in Section 5 with the current optimal approach based on the Benes network (Gentry et al., 2012; Halevi and Shoup, 2014). The average performance on randomly generated permutations with different lengths $n$ is presented in the Table. The benchmarks are constructed as follows: both schemes are initially set with a multiplication depth of $\log{n}-1$ (aligned with the configuration in Table 2). Then, each scheme is applied with its level-collapsing strategy to measure the optimal performance. During the tests, computations with the same $n$ all end at modulus level $117-\log{n}+$ 1.
The Benes network-based solution employs the level-collapsing strategy proposed by Halevi et al.(Halevi and Shoup, 2014). On the other hand, the best practice for our network is to compress the top $2\sim 3$ levels and the bottom $3\sim 4$ levels. Compressing these levels reduces both the multiplication depth and the number of rotations (recall Table 2). Rotations for the compressed levels are further accelerated with the hoisting technique.
The Speed-up column in Table 7 reveals that our algorithm provides speed increase across different $n$. Given the typical setup range of $10\leq\log{N}\leq 16$ for RLWE-based batch-encoding homomorphic encryption schemes (Bossuat et al., 2024), our algorithm demonstrates an efficiency advantage over nearly all practical choices of $n$.
@@ Appendix A Batch-encoding HE Settings
  • $\texttt{Setparams}(\cdot)$: Generate parameters $params$: $(N,Q,P,\chi_{sec},\chi_{err})$. $N$ is a power of two, which along with $Q$ and $P$ defines the polynomial rings $R=\mathbb{Z}[X]/(X^{N}+1)$, $R_{Q}=\mathbb{Z}_{Q}[X]/(X^{N}+1)$, and $R_{QP}=\mathbb{Z}_{QP}[X]/(X^{N}+1)$, where $Q=\prod_{i=0}^{L-1}q_{i}$, $P=\prod_{i=0}^{\alpha-1}p_{i}$, and any $q_{i}$ for $0\leq i<L$ or $p_{i}$ for $0\leq i<\alpha$ are congruent to $1$ modulo $2N$. The distributions $\chi_{sec}$ and $\chi_{err}$, defined over $R$, are used for sampling the secret key and noise, respectively. All subsequent routines implicitly receive $params$ as input.
@@
  • $\texttt{Enc}(m,\ell)$: For the given plaintext $m\in R$ and the index $\ell$ of the modulus chain $Q$, compute $(u,v)\in R_{Q_{\ell}}^{2}\leftarrow(ar+e_{2},(-as+e_{1})r+\Delta_{1}m+\Delta_{2}e_{3})$, where $Q_{\ell}=\prod_{i=0}^{\ell-1}q_{i}$ and $e_{1},e_{2},e_{3}\leftarrow\chi_{err}$. $\Delta_{1}$ and $\Delta_{2}$ are scaling factors used to control the precision of the plaintext, where one of them must be $1$, while the other scales the polynomial coefficients, covering both low-bit and high-bit encoding cases for $m$.
@@ B.1. Partition pattern of arbitary dimension
Ours (Table 4) $d/1$ 512 7 ms 9 ×$\times 1.00$ 128 72 ms 12 ×$\times 1.00$ 32 694 ms 15 ×$\times 1.00$
Ours (Fast Rep) $d/2$ 512 5 ms 9 ×$\times 1.33$ 128 56 ms 12 ×$\times 1.28$ 32 567 ms 15 ×$\times 1.22$
Ours (Fast Rep) $d/4$ 512 4 ms 12 ×$\times 1.76$ 128 42 ms 15 ×$\times 1.72$ 32 471 ms 18 ×$\times 1.47$
Ours (Fast Rep) $d/8$ 512 2 ms 21 ×$\times 3.13$ 128 30 ms 24 ×$\times 2.35$ 32 354 ms 27 ×$\times 1.96$
Ours (Fast Rep) $d/16$ 128 21 ms 45 ×$\times 3.40$ 32 279 ms 48 ×$\times 2.49$
Ours (Fast Rep) $d/32$ 32 232 ms 93 ×$\times 2.99$
@@ B.2. Demonstration of $1$-padding on $U^{\gamma}$ and $U^{\xi}$
For any $d^{2}\times d$ matrices $\hat{A}$ and $\hat{B}$, $U^{\gamma}_{pad}(\hat{A})$ has the same first column as $U^{\gamma}(\hat{A})$, and $U^{\xi}_{pad}(\hat{B})$ has the same first $d$ columns as $U^{\xi}(\hat{B})$.
@@
For $U^{\gamma}(\hat{A})$, entries of $A$ in $\hat{A}$ are rotated by $-k(d^{2}-1)$ steps for some $0\leq k<d$ to ensure those with column index $i$ in equivalence class $i\bmod{d}$ are shifted to equivalence class $0\bmod{d}$. Since $(d^{2}-1)$ is coprime to $d$, for each equivalence class $i\bmod{d}$, there is exactly one unique $k$ such that $i+k=0\bmod{d}$. This implies that during each rotation by steps $0,-(d^{2}-1),\dots,-(d^{2}-1)(d-1)$, only one equivalence class is rotated to the first column position of $\hat{A}$. The correctness of $U^{\xi}_{pad}(\hat{B})$ is consistent with the discussion for $U^{\gamma}$ above. $U^{\xi}(\hat{B})$ requires moving $\mathbf{b}_{i}$ in $\hat{B}$ by $k(d-1)$ positions (with $1\times d$ vectors as units) to the column indexed by 0, where $0\leq i,k<d$. For each $\mathbf{b}_{i}$, there is exactly one unique $k$ such that $i+k(d-1)=0\mod d$. This implies that during the rotations of $\hat{B}$ by steps of $0,-(d-1),\dots,-(d-1)^{2}$, each rotation carries exactly one $\mathbf{b}_{i}$ to the first column position. ∎
@@ B.3. Optimization of Replication
This routine requires $O(d\log{d})$ rotations to produce replications of all elements in $M$. To optimize this, we divide $d=d_{1}d_{0}$ and perform rotations with steps not smaller than $d_{0}$ before masking:
@@
where $\mathbf{u}_{k,i}$ extracts $\mathbf{m}_{k}$ from $\texttt{Rot}\left(M;d_{0}i\right)$. The pre-rotated components $\texttt{Rot}(M;d_{0}i)$ can be cached for all $d$ elements’ computation but only costs $d_{1}$ rotations. The total number of rotations is $\left(\frac{d}{d_{0}}+d\log{d_{0}}\right)$, and the multiplication depth remains $1$. For HMM regarding $M$ as a matrix, this method can directly apply to row-wise replication. Extending the pre-rotated steps to {±$\{\pm d_{0}i\mid 0\leq i<d_{1}\}$ enables column-wise replication with a number of $\left(\frac{2d}{d_{0}}+d\log d_{0}\right)$ rotations. Therefore, the number of rotations for our HMM in Section 4.2.2 is reduced to $\left(\frac{3d}{md_{0}}+\frac{2d}{m}\log{d_{0}}\right)$ for $d^{\prime}=1$, surpassing a range of schemes with the complexity of $O(d\log d)$ under $d^{2}$ encoding space (Cheon et al., 2018; Huang and Zong, 2023; Aikata and Roy, 2024). The corresponding performance, along with a comparison to the baseline version without this optimization, is presented in Table 8.

### v7

@@ 6.1. Homomorphic Matrix Transposition
Table 3 provides HMT benchmarks under dimension $d\in\{128,256\}$ and decomposition depth $\ell\leq\lfloor\log{(d-1)}\rfloor$. Both the depth-$\ell$ decomposed HMT and the original version used for comparison terminate at modulus level $17-(\ell+1)$. To support HMT of $d=256$, the polynomial degree modulus $N$ is extended to $2^{16}$, and the encoding capacity is alternated to $N$ real numbers by the conjugate-invariant ring based CKKS (Kim and Song, 2019). As previously discussed, the leftmost factor $U^{t}_{L_{\ell}}$ is implemented with the BSGS algorithm. $U^{t}_{L_{\ell}}$ contains $2d/2^{\ell}$ non-zero diagonals, distributed symmetrically on both sides of the 0-th diagonal. Thus, the product of BSGS loop sizes $d_{1}d_{2}$ is no greater than $d/2^{\ell}$, the number of non-zero diagonals on one side of $U^{t}_{L_{\ell}}$.
Compared to the original undecomposed version (Jiang et al., 2018), the ideal decomposition yields significant complexity reductions, even when the decomposition depth is shallow. This aligns with the reduced number of rotations from $d_{1}+\frac{2d}{d_{1}}$ to $2\ell+d_{1}+\frac{2d}{2^{\ell}d_{1}}$ by depth-$\ell$ decomposition. Notably, our method demonstrates a speed-up of up to $2.3\times$ and a rotation key reduction of up to $67\%$ for $d=256$. The asymptotic complexity improvement can also be evidenced by the runtime growth with respect to $d$. While the original $U^{t}$ exhibits a consistent growth factor of $3.7$, the decomposed version shows a decreasing trend from $3.3$ to $2.6$ as $\ell$ increases.
@@ 6.2. Homomorphic Matrix Multiplication
Table 4 primarily compares our enhanced HMM in Section 4.2 to the original scheme (Rizomiliotis and Triakosia, 2022). Unlike our variable $d^{2}d^{\prime}$ matrix encoding space, the original $d^{3}$-encoding only allows limited choices of dimensions. Thus, the ciphertext’s encoding capacity is also shifted to $N$ real numbers (Kim and Song, 2019) (HE parameters remain default). All computations terminate at modulus level $14$.
In the table, parameter $m=N/d^{2}d^{\prime}$ denotes the maximum number of matrices that can be computed in parallel. For the original encoding setting $d^{\prime}=d$, our scheme achieves a speedup of up to $3.24\times$ and a 54% reduction in rotation keys compared to the original algorithm. This improvement is purely achieved by the permutation decomposition technique, which allows the complexity to grow logarithmically with $d$ and remain independent of multiplication depth overhead.
The rest of the experiments under different $d^{\prime}$ demonstrate the enhancement of our more flexible encoding space utilization method. We can trade encoding space for higher parallelity and lower amortized time per matrix multiplication. The HMM algorithm proposed by Jiang et al. is also included for reference (Jiang et al., 2018). Here, we adopt the version enhanced by Ma et al. through hoisting techniques (Ma et al., 2024). Recall Table 1, this HMM algorithm achieves optimal amortized complexity. However, its complexity is fixed over different batch sizes of matrix multiplication. Thus, when it comes to sequential and small-batch matrix multiplications, our approach offers a solution with significantly lower latency. A representative scenario of such use cases is explored in the following subsection.
@@
[…]
@@
    Original implementation under Jiang et al.’s HMM (Jiang et al., 2018; Ma et al., 2024).
@@
$\log{n}$ Time (ms) Collapsed Time (ms) Collapsed Speed-up
77 308 3 187 Top 0 Bot 2 $\times 1.65$
88 426 3 252 Top 0 Bot 2 $\times 1.69$
99 553 5 362 Top 0 Bot 3 $\times 1.53$
$10$ 655 4 481 Top 0 Bot 3 $\times 1.36$
$11$ 834 5 617 Top 2 Bot 3 $\times 1.35$
$12$ 1013 5 810 Top 2 Bot 3 $\times 1.25$
$13$ 1184 4 1030 Top 2 Bot 3 $\times 1.15$
$14$ 1405 6 1244 Top 2 Bot 3 $\times 1.13$
@@
6.3. Application to Oblivious Neural Network Inference
Oblivious inference refers to the task of outsourcing an ML model without revealing its parameters or inputs, wherein the model owner is not required to run the model locally (Jiang et al., 2018; Chen et al., 2019; Rizomiliotis et al., 2022). Jiang et al. (Jiang et al., 2018) proposed an efficient HE-based framework, E2DM, to realize this objective, predicting the MNIST dataset (Deng, 2012) over an encrypted neural network. The computational complexity of this framework is dominated by their proposed HMM. We adopt their service paradigm but replace the HMM module in their framework with our version. Our objective is to demonstrate the practicality of our HMM by achieving lower latency when redundant encoding capacity is available, particularly under small-batch inference scenarios.
Table 5 describes the employed neural network. We call it E2DM-lite, as it is derived by halving the number of convolutional channels and neurons in E2DM, but achieves an accuracy of 97.7% (only 0.4% lower than the original version). We present benchmarks of E2DM-lite in Table 6 (implementation detail provided in Appendix E). When implemented with the original HMM by Jiang et al., E2DM-lite requires an inference batch size of $128$ for optimal performance. Notably, its latency remains unchanged when the batch size drops to $32$, as the framework utilizes parallel HMMs of size $32\times 32$ to process fully connected layers. When the batch size further decreases to $16$ or $8$, the parallel HMM blocks become smaller, and the cost of aggregating these blocks after HMM increases correspondingly. In contrast, our HMM takes advantage of the encoding space made available by small batches, achieving latency improvements of $3.0\times$, $3.5\times$, and $3.9\times$ for batch sizes of 32, 16, and 8, respectively, while maintaining an amortized runtime that is only slightly higher than that of (Jiang et al., 2018). Note that although it is possible to implement the original E2DM with our proposed techniques, the HMM size in E2DM is defined relatively larger, where using a redundant encoding space for latency reduction would result in increased polynomial degree modulus $N$, posing a negative effect on the overall performance of homomorphic operations.
Table 6 also presents benchmarks of other state-of-the-art HE-based oblivious neural network inference frameworks on the MNIST dataset. Note that although they use similar underlying HE schemes (Cheon et al., 2017, 2019) and network topologies, we cannot force them to work under identical parameters of HE and neural network, which have been highly customized for different computation strategies. Therefore, the benchmarks are obtained under the original parameters of these frameworks, except for E2DM, which only ensures 80-bit security. We have it implemented under the same HE setting as our scheme.
Jiang et al.’s HMM implementation in E2DM and E2DM-lite exhibits good amortized performance under high batch sizes, but introduces relatively high depth, making it more suitable for shallow, wide networks. The strategy designed by Chen et al.(Chen et al., 2019) (CDKS) employs their proposed multi-key homomorphic encryption to avoid introducing a third-party authority. It has a good latency, but the batch size is relatively small. POSEIDON(Sav et al., 2021) is a federated neural network learning framework based on threshold HE. It provides fast oblivious inference after the network is homomorphically trained among parties. However, clients outside the training system need to encrypt their inquiries with the joint public key of the system.
@@ 6.4. Arbitrary Homomorphic Permutation
Table 7 compares our proposed homomorphic permutation algorithm in Section 5 with the current optimal approach based on the Benes network (Gentry et al., 2012; Halevi and Shoup, 2014). The table reports the average performance on randomly generated permutations of various lengths $n$. The homomorphic evaluation for all permutation lengths is configured to terminate at modulus level 0. Both schemes initially constructs networks with $\log{n}-1$ multiplication depth, then apply their respective level-collapsing strategies to seek optimal performance, where the reduced number of network levels is reported in the table.
The Benes network-based solution employs the level-collapsing strategy proposed by Halevi et al.(Halevi and Shoup, 2014). For our network, the best practice is to compress the top $2\sim 3$ levels and the bottom $3\sim 4$ levels. Compressing these levels reduces the number of homomorphic operations performed on relatively high modulus levels. Rotations for the compressed levels are further accelerated with the hoisting technique. The Speed-up column in Table 7 reveals that our algorithm provides speed increase across different $n$. Given the typical setup range of $10\leq\log{N}\leq 16$ and $n\leq N$ for RLWE-based batch-encoding homomorphic encryption schemes (Bossuat et al., 2024), our algorithm demonstrates an efficiency advantage over nearly all practical choices of $n$.
@@ Appendix A Batch-encoding HE Settings
  • $\texttt{Setparams}(\cdot)$: Generate parameters paramsparams: $(N,Q,P,\chi_{sec},\chi_{err})$. $N$ is a power of two, which along with $Q$ and $P$ defines the polynomial rings $R=\mathbb{Z}[X]/(X^{N}+1)$, $R_{Q}=\mathbb{Z}_{Q}[X]/(X^{N}+1)$, and $R_{QP}=\mathbb{Z}_{QP}[X]/(X^{N}+1)$, where $Q=\prod_{i=0}^{L-1}q_{i}$, $P=\prod_{i=0}^{\alpha-1}p_{i}$, and any $q_{i}$ for $0\leq i<L$ or $p_{i}$ for $0\leq i<\alpha$ are congruent to $1$ modulo $2N$. The distributions $\chi_{sec}$ and $\chi_{err}$, defined over $R$, are used for sampling the secret key and noise, respectively. All subsequent routines implicitly receive paramsparams as input.
@@
  • $\texttt{Enc}(m,\ell)$: For the given plaintext $m\in R$ and the index $\ell$ of the modulus chain $Q$, compute $(u,v)\in R_{Q_{\ell}}^{2}\leftarrow(ar+e_{2},(-as+e_{1})r+\Delta_{1}m+\Delta_{2}e_{3})$, where $Q_{\ell}=\prod_{i=0}^{\ell-1}q_{i}$ and $e_{1},e_{2},e_{3}\leftarrow\chi_{err}$. $\Delta_{1}$ and $\Delta_{2}$ are scaling factors used to control the precision of the plaintext, where one of them must be 11, while the other scales the polynomial coefficients, covering both low-bit and high-bit encoding cases for $m$.
@@ B.1. Partition pattern of arbitary dimension
Ours (Original) $d/1$ 512 7 ms 9 $\times 1.00$ 128 72 ms 12 $\times 1.00$ 32 694 ms 15 $\times 1.00$
Ours (Fast Rep) $d/2$ 512 5 ms 9 $\times 1.33$ 128 56 ms 12 $\times 1.28$ 32 567 ms 15 $\times 1.22$
Ours (Fast Rep) $d/4$ 512 4 ms 12 $\times 1.76$ 128 42 ms 15 $\times 1.72$ 32 471 ms 18 $\times 1.47$
Ours (Fast Rep) $d/8$ 512 2 ms 21 $\times 3.13$ 128 30 ms 24 $\times 2.35$ 32 354 ms 27 $\times 1.96$
Ours (Fast Rep) $d/16$ - - - - 128 21 ms 45 $\times 3.40$ 32 279 ms 48 $\times 2.49$
Ours (Fast Rep) $d/32$ - - - - - - - - 32 232 ms 93 $\times 2.99$
@@ B.2. Demonstration of $1$-padding on $U^{\gamma}$ and $U^{\xi}$
[…]
@@
[…]
Theorem 2 demonstrates the correctness of the $1$-padding version of $U^{\gamma}$ and $U^{\xi}$. We now focus on the $1$-padding of their respective decompositions. Following the decomposition pattern of $U^{t}$, the decomposition of $U^{\gamma}$ is typically given as:
[…]
where ”*” indicates the block containing zeros and meaningless values. Similarly, for $1\leq\ell<\log{d}$, a $U^{\gamma}_{pad_{R_{\ell}}}$ rotates its input with distance $(d^{2}-1)\cdot d/2^{\ell}\mod{d^{3}}$ and adds the rotated result to the original input. The behaviour of such $U^{\gamma}_{pad_{R_{\ell}}}$ is interpreted as follows: it evenly partitions each block of the input matrix into $4$ subblocks and transposes the top two subblocks within each leftmost block into the first column, while the remaining subblocks are left with either zeros or garbage values. Note that the initial block is the input matrix $\hat{A}$ itself. Then, $U^{\gamma}_{pad_{L_{\ell}}}$ is defined as performing $U^{\gamma}_{L}$ on each leftmost $d^{2}/2^{\ell}\times d/2^{\ell}$ block of its input while zeroing out all other positions. This operation costs $1$ multiplication depth.
The $1$-padding version of the decomposed $U^{\xi}$ is defined exactly the same way as $U^{\gamma}$, except that the size of the input matrix is $d\times d^{2}$, whereas the size of the input matrix of $U^{\gamma}$ is considered as $d^{2}\times d$.
@@ B.3. Optimization of Replication
This method requires $O(t\log d)$ rotations to replicate $t$ elements. To improve efficiency, we decompose $d=d_{1}d_{0}$ and perform rotations with step sizes at least $d_{0}$ prior to masking:
@@
where $\mathbf{u}_{k,i}$ extracts $\mathbf{m}_{k}$ from $\texttt{Rot}\left(M;d_{0}i\right)$. The pre-rotated components $\texttt{Rot}(M;d_{0}i)$ can be cached for all $d$ elements’ computation but only costs $d_{1}$ rotations. The total number of rotations is $\left(\frac{d}{d_{0}}+d\log{d_{0}}\right)$, and the multiplication depth remains $1$. For HMM regarding $M$ as a matrix, this method can directly apply to row-wise replication. Extending the pre-rotated steps to $\{\pm d_{0}i\mid 0\leq i<d_{1}\}$ enables column-wise replication with a number of $\left(\frac{2d}{d_{0}}+d\log d_{0}\right)$ rotations. Therefore, the number of rotations for our HMM in Section 4.2.2 is reduced to $\left(\frac{3d}{md_{0}}+\frac{2d}{m}\log{d_{0}}\right)$ for $d^{\prime}=1$. From an asymptotic perspective, this rotation complexity improves upon a range of schemes with $O(d\log d)$ complexity under $d^{2}$ encoding space (Cheon et al., 2018; Huang and Zong, 2023; Aikata and Roy, 2024). The corresponding performance, along with a comparison to the baseline version without this optimization, is presented in Table 8.